Janet is three times as old as her daughter, Mary.Five years ago, Janetwas 4 times as old as Mary. How old is Janet now? How old willbe Mary in 7 years time?
Answer by jorel1380(3719)

j=3m
j-5=4(m-5)
3m-5=4m-20
m=15
j=45
Janet is 45. Mary will be 22 in 7 years..